Yunikon:

The sum of two numbers is 8. Is it possible for the difference of the two numbers to also be 8? Why or why not?

darkknight:

a+b = 8 a-b or b-a = 8?

darkknight:

try plugging numbers in, such as 8 and 0 for a and b

Yunikon:

i have to explain if its possible or not

xxDeppressionxx:

Yes, it is possible for two numbers to have a difference of 8 and a sum of 1.

darkknight:

two numbers: a and b sum is 8 So a + b = 8 now a - b or b - a = 8? plug in numbers for a and b and see if they satisfy the second statement

Yunikon:

Thx ^﹏^

jhonyy9:

- for example suppose a +b = 8 => a = 8 - b in case of a -b = 8 than you substitute the 8 - b in place of a you get 8 - b - b = 8 8 - 2b = 8 - 2b = 0 b = 0 and now than substitute this b value of 0 in place of b inside this equation a +b = 8 so a + 0 = 8 => a = 8 so for a = 8 and b = 0 it's possible or invers a = 0 and b = 8 hope helped easy understandably